OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 4)(2x – 1)(x – 7)?

OpenStudy (zarkon):

solve x + 4=0 2x – 1=0 x – 7=0

OpenStudy (anonymous):

oooh ok thanks i got it